OpenStudy (anonymous):

Approximately 85% of applicants get their G1 drivers license the first time trying the test. if 80 applicants take the test, what is the probability more than 10 applicants will have to retake the test. Use normal approximation to binomial distribution

OpenStudy (perl):

so we want P( X >= 10 )

OpenStudy (perl):

the probability of passing the test the first time is .85. the probability of having to retake the test is 1-.85

OpenStudy (perl):

ie failing the test

OpenStudy (anonymous):

is 0.15

OpenStudy (perl):

right :)

OpenStudy (anonymous):

how would I get the z-score?

OpenStudy (perl):

you would have to use a binomial continuity correction , as well

OpenStudy (anonymous):

The mean of the binomial distribution is \(np\), where \(n\) is the number of trials and \(p\) is probability of success.

OpenStudy (anonymous):

so 10.5?

OpenStudy (perl):

80*.15 = 12

OpenStudy (anonymous):

While the standard deviation is given by \(\sqrt{npq}\) where \(q=1-p\) or probability of failure.

OpenStudy (perl):

teh standard deviation is sqrt( n * p * (1-p))

OpenStudy (perl):

the*

OpenStudy (anonymous):

when finding the mean, np, is p .85 or .15

OpenStudy (anonymous):

So our z score will be: \[ z = \frac{x-np}{\sqrt{np(1-p)}} \]where \(x\) is the number of success we want

OpenStudy (anonymous):

In our case, success would mean that you have to retake the test, ironically

OpenStudy (perl):

You are going to want P( X >= 9.5) because you are approximating a discrete distribution with a continuous one. alternatively you can also solve this problem with a complement

OpenStudy (anonymous):

okay I got it, thank you :)

OpenStudy (briellshaw):

1/8 as a fraction it would be 1 out of 8 i believe i don't know much i'm only in the seventh grade